Please help me solve : x(5x+17)=12
Found 2 solutions by ewatrrr, amalm06:
Answer by ewatrrr(24785)


Hi
Solving:
x(5x+17)=12
5x^2 + 17x = 12
5x^2 + 17x - 12 = 0
factoring
(5x -3)(x +4)= 0 Note:SUM of the inner product(-3x) and the outer product(20x) = 17x
(5x -3)=0
x = 3/5
(x +4)= 0
x = -4
CHECKING our Answer(s)***
3/5(3+17 = 12
-4(-3) = 12

Answer by amalm06(224)

Write the equation in quadratic form and then use the quadratic formula to find the real solutions. Or simply factor.